When lithium metal is irradiated with light with λ = 300 nm a stopping potential of 1.83 V is observed. Calculate the threshold frequency and work function for lithium.


1
Expert's answer
2021-02-09T03:46:42-0500

Potential = 1.83 V

for one electronic charge

one coulomb×electronic charge

Potential = 1.83eV

E = "\frac{hc}{\lambda}"

E = "\frac{6.626×10^-³⁴ × (3×10⁸)}{300×10^-⁹}"

E = 6.626×10-19

Now , hv = (6.626×10-19)/( 1.6×10-19 )

hv = 4.141

E = 4 141-1.83 = 2.311 eV

Now in Joule

E = 2.311×1.6 × 10-19

E = 3.6976 × 10-19

Now threshold frequency

"\nu" =E/h = (3.6976×10-19)/(6.626×10-34)

"\nu" = 55.8 × 1013 sec-1